OpenStudy (anonymous):

Solve for x: 7 over 8 minus 1 over x equals 3 over 4 x = 1 x = 2 x = 4 x = 8

OpenStudy (anonymous):

7/8 - 3/4 = 1/8 1/? ? = 8

OpenStudy (anonymous):

So it is D.

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

Thank you so much! thats what I was thinking

OpenStudy (anonymous):

It does!

OpenStudy (anonymous):

Thats Awesome!